In a batch of 8,000 clock radios 4% are defective. A sample of 6 clock radios is randomly selected without replacement from the 8,000 and tested. The entire batch will be rejected if at least one of those tested is defective. What is the probability that the entire batch will be rejected?

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

In a batch of 8,000 clock radios 4% are defective. A sample of 6 clock radios is randomly selected without replacement from the 8,000 and tested. The entire batch will be rejected if at least one of those tested is defective. What is the probability that the entire batch will be rejected?
------------------------
Binomial Problem with n=6 and p = 0.04
----
P(at least one defective) = 1 - P(none are defective)
---
= 1 - 0.96^6
---
= 1 - 0.7828
---
= 0.2172
==============
Cheers,
Stan H.
==============

Answer by edjones(8007)

8000*.04=320
8000-320=7680
7680C6/8000C6 = .7827 probability none found defective.
1-.7827=.2173 the probability that the entire batch will be rejected.
.
Ed